# Injecting Hierarchical Biological Priors into Graph Neural Networks for Flow Cytometry Prediction

### v2

@@ 1 Introduction
Flow cytometry (FC) is a technology that provides rapid multi-parametric analysis of single cells in solution. The expression of specific cell surface and intracellular molecules is thereby detected by antibodies coupled to fluorochromes. Exposing labelled cells to a laser leads to the emission of fluorescence which is then collected by a detector. By analyzing more than millions of cells in a short time, information can thereby be collected about the presence or absence of up to fifty different cellular molecules per cell. As a result, histologists obtain tabular data where each row corresponds to a single cell and every column to a marker or feature (cell surface molecule – Table 4, in Appendix A).
@@
In this work, we propose a novel approach to inject hierarchically-structured priors (biological in our case) into the GNN framework for multi-class prediction tasks on tabular data. Specifically, we encode the known hierarchical relationships between different cell types or functional classes as a tree-structured hierarchy imposed on the GNN output space. This hierarchical constraint serves as an inductive bias that encourages the GNN to respect the hierarchical dependencies inherent to the biological domain. During training, we use a custom hierarchical loss function that accounts for the hierarchical similarities between classes, in addition to the traditional cross-entropy loss. This approach ensures that the model’s predictions not only accurately classify instances into their respective leaf nodes (specific cell types), but also respect the hierarchical groupings at higher levels of the taxonomy (broader cell lineages or functional categories). Through rigorous experiments on our real-world datasets, we demonstrate that incorporating hierarchical biological priors significantly boosts the performance of GNN models across a range of metrics. Our hierarchical GNN approach achieves consistent gains over strong baselines that do not leverage such structured domain knowledge. We provide in-depth analysis and empirical evidence highlighting the importance of encoding inductive biases aligned with the underlying data domains, especially for domains with rich hierarchical relationships like cell-biology.
@@ 2 Related Work
The primary research focus in machine learning has predominantly been centered on developing models for conventional classification problems, wherein an object is assigned to a single class from a set of disjoint classes. However, a distinct subset of tasks involves scenarios where classes are not disjoint but rather organized hierarchically, giving rise to hierarchical classification (HC). The hierarchical structure formalizing the interrelation among classes can manifest as either a tree or a directed acyclic graph (DAG) (Silla & Freitas, 2011). Hierarchical classification problems manifest across a broad spectrum, spanning from musical genre classification (Ariyaratne & Zhang, 2012; Iloga et al., 2018) to identification of COVID-19 in chest X-ray images (Pereira et al., 2020), taxonomic classification of viral sequences in metagenomic data (Shang & Sun, 2021) and text categorization (Javed et al., 2021; Ma et al., 2022). Graph-based models have the capability to leverage both the global and local characteristics inherent in networks relevant to cell biology. GNNs have been recently adapted in various tasks, e.g. link prediction (Zhang & Chen, 2018), graph classification (Duvenaud et al., 2015; Lee et al., 2019), and node classification (Kipf & Welling, 2017; Velickovic et al., 2017; Yang et al., 2016). In the context of HC problems, DiffPool (Ying et al., 2018) propose a differentiable graph pooling module to generate hierarchical representations of graphs using various GNNs in an end-to-end fashion. Moreover, in the biological domains, HC-GNNs have been widely applied for histological image classification (Hou et al., 2022), protein-protein interaction (Gao et al., 2023), molecules’ properties prediction (Han et al., 2023) and for cancer diagnosis and prognogis in digital pathology (Pati et al., 2022).
@@ 6 Conclusion and Future Work
We have presented a novel hierarchical GNN framework, namely FCHC-GNN, that injects structured biological priors to enhance multi-class prediction on FC cellular data. By encoding known hierarchical relationships between cell types and functional classes, our model effectively captures the rich dependencies innate to biological domains while operating on a tabular input representation. Extensive experiments across ours FC dataset demonstrate substantially improved performance over strong baselines that fail to leverage such hierarchical constraints. Moreover, as demonstrated over well-know ImageNet data, our FCHC-GNN module is easily extendible to other domain where classes may be organized within an hierarchy.
The consistent gains validate our hypothesis that tailored inductive biases aligned with the underlying tabular data distributions are crucial for achieving better generalization. FCHC-GNN architectures introduces a novel method to enhance learning on tabular inputs with prior domain knowledge expressed as hierarchies. Looking ahead, this general approach holds promise for biomedical learning tasks beyond single cell classification.
@@ Appendix D ImageNet1000-Mini Results
Table 5 shows competitive performance of our FCHC-GNN module on hierarchical image classification task as well. For sake of simplicity, we only showed our plug-in module being applied to a GAT, against its normal version and the average state-of-art performances (accordingly to what written on the benchmark https://paperswithcode.com/sota/image-classification-on-imagenet).

### v3

@@ 1 Introduction
Flow cytometry (FC) is a technology that provides rapid multi-parametric analysis of single cells in solution. The expression of specific cell surface and intracellular molecules is thereby detected by antibodies coupled to fluorochromes. Exposing labeled cells to a laser leads to the emission of fluorescence which is then collected by a detector. By analyzing more than millions of cells in a short time, information can thereby be collected about the presence or absence of up to fifty different cellular molecules per cell. As a result, histologists obtain tabular data where each row corresponds to a single cell and every column to a marker or feature (cell surface molecule – Table 4, in Appendix A).
@@
In this work, we propose a novel approach to inject hierarchically-structured priors (biological in our case) into the GNN framework for multi-class prediction tasks on tabular data. Specifically, we encode the known hierarchical relationships between different cell types or functional classes as a tree-structured hierarchy imposed on the GNN output space. This hierarchical constraint serves as an inductive bias that encourages the GNN to respect the hierarchical dependencies inherent to the biological domain. During training, we use a custom hierarchical loss function that accounts for the hierarchical similarities between classes, in addition to the traditional cross-entropy loss. This approach ensures that the model’s predictions not only accurately classify instances into their respective leaf nodes (specific cell types), but also respect the hierarchical groupings at higher levels of the taxonomy (broader cell lineages or functional categories). Through rigorous experiments on our real-world datasets, we demonstrate that incorporating hierarchical biological priors significantly boosts the performance of GNN models across a range of metrics. Our hierarchical GNN approach achieves consistent gains over strong baselines that do not leverage such structured domain knowledge. We provide in-depth analysis and empirical evidence highlighting the importance of encoding inductive biases aligned with the underlying data domains, especially for domains with rich hierarchical relationships like cell biology.
@@ 2 Related Work
The primary research focus in machine learning has predominantly been centered on developing models for conventional classification problems, wherein an object is assigned to a single class from a set of disjoint classes. However, a distinct subset of tasks involves scenarios where classes are not disjoint but rather organized hierarchically, giving rise to hierarchical classification (HC). The hierarchical structure formalizing the interrelation among classes can manifest as either a tree or a directed acyclic graph (DAG) (Silla & Freitas, 2011). Hierarchical classification problems manifest across a broad spectrum, spanning from musical genre classification (Ariyaratne & Zhang, 2012; Iloga et al., 2018) to the identification of COVID-19 in chest X-ray images (Pereira et al., 2020), the taxonomic classification of viral sequences in metagenomic data (Shang & Sun, 2021) and text categorization (Javed et al., 2021; Ma et al., 2022). Graph-based models can leverage both the global and local characteristics inherent in networks relevant to cell biology. GNNs have been recently adapted in various tasks, e.g. link prediction (Zhang & Chen, 2018), graph classification (Duvenaud et al., 2015; Lee et al., 2019), and node classification (Kipf & Welling, 2017; Velickovic et al., 2017; Yang et al., 2016). For FC data FlowCyt (Bini et al., 2024) represents the first publicly available benchmark, designed to test deep learning models for single-cell multi-class classification, highlighting the benefits of modeling these data as a graph-structured problem. In the context of HC problems, DiffPool (Ying et al., 2018) propose a differentiable graph pooling module to generate hierarchical representations of graphs using various GNNs in an end-to-end fashion. Moreover, in the biological domains, HC-GNNs have been widely applied for histological image classification (Hou et al., 2022), protein-protein interaction (Gao et al., 2023), molecules’ properties prediction (Han et al., 2023) and for cancer diagnosis and prognosis in digital pathology (Pati et al., 2022).
@@ 6 Conclusion and Future Work
We have presented a novel hierarchical GNN framework, namely FCHC-GNN, that injects structured biological priors to enhance multi-class prediction on FC cellular data. By encoding known hierarchical relationships between cell types and functional classes, our model effectively captures the rich dependencies innate to biological domains while operating on a tabular input representation. Extensive experiments across our FC dataset demonstrate substantially improved performance over strong baselines that fail to leverage such hierarchical constraints. Moreover, as demonstrated over well-known ImageNet data, our FCHC-GNN module is easily extendible to other domains where classes may be organized within a hierarchy.
The consistent gains validate our hypothesis that tailored inductive biases aligned with the underlying tabular data distributions are crucial for achieving better generalization. FCHC-GNN architectures introduce a novel method to enhance learning on tabular inputs with prior domain knowledge expressed as hierarchies. Looking ahead, this general approach holds promise for biomedical learning tasks beyond single-cell classification.
@@ Appendix D ImageNet1000-Mini Results
Table 5 shows a competitive performance of our FCHC-GNN module on hierarchical image classification task as well. For the sake of simplicity, we only showed our plug-in module being applied to a GAT, against its normal version and the average state-of-art performances (according to what was written on the benchmark https://paperswithcode.com/sota/image-classification-on-imagenet).